### v1

# Closure for Microscale Fluctuations in Chaotic Mixing Fronts

Abstract
Scalar mixing fronts are formed under the stirring and spreading action of a random velocity field. The prediction of concentration statistics below the velocity length scale is an open problem. Here, we show that when fluid agitation is caused by a smooth chaotic flow with weak persistence, the evolution of scalar variance can be uniquely related to the stirring and spreading properties of the flow and the diffusivity of the scalar. Using the spectral description of microscale scalar fluctuations in presence of a large-scale forcing by Kraichnan [1], we uncover the link between variance production and dissipation in a dispersing front. This provides a universal closure for the scalar fluctuation microscale whenever there exists a clear separation between macroscale dispersive forcing and microscale dissipation. We validate the closure model on high-resolution direct numerical simulation of advection-diffusion in chaotic flows of variable strength and persistence. These findings open a new avenue to predict both conservative and reactive transport in incompletely mixed environments, ranging from porous media flows to engineered flows in microfluidic devices.
preprint: APS/123-QED
I Introduction
Figure 1: Scalar front mixed and dispersed by a smooth periodic chaotic flow (sine flow). Scalar concentrations (colorscale) obey Eq. (7) with $\kappa=5\cdot 10^{-6}$ and $\bm{u}$ defined by Eq.(LABEL:eq:sineflow), with $U=0.2$ ($\gamma\approx\sigma^{2}_{\gamma}\approx 0.05$), and $t=300$. All the important length scales appearing in the front mixing problem are shown at real scale in the plot (from small to large). For the considered scenario, the Batchelor scale is $s_{B}\approx 0.01s_{v}$  (1), the modified Batchelor scale is $s_{B}^{\prime}\approx 0.01s_{v}$ (Eq. (28)), the microscale is $s_{m}\approx 0.02s_{v}$ (Eq. (19)), the velocity scale is $s_{v}\equiv 1$, the injection scale is $s_{i}\approx 2.7s_{v}$ (Eq. (36)) and the dispersive scale $s_{d}\approx 1.41s_{v}$ (Eq. (4)).
Solute mixing by chaotic fluid flows occurs in a range of natural and industrial processes, including static mixers [2], turbulence  [3], or porous media [4]. Away from the source, mixing fronts naturally develop between the diluting and dispersing solute and the surrounding fluid, forming concentration gradients over a wide range of length scales (Fig. 1). Such interfaces are observed, for instance, in the fringe of contaminant plumes in the subsurface [5], in density or thermal currents in lakes or coastal areas, or in river junctions [6]. A universal modelling framework that captures the mixing process in such interfaces is still missing, even in the simple case where the solute is passively transported by the flow and the flow is smooth at the scale of interest. In contrast to rough turbulent flows, smooth chaotic flows are characterized by a minimum velocity lengthscale $s_{v}$ below which velocity gradients are constant (Fig. 1). Such conditions are found for laminar flows at low Reynolds numbers or in turbulent flows at high Schmidt numbers, where dissipative scales are larger than mixing scales. Below $s_{v}$, solute mixing occurs by the simultaneous action of fluid deformation and molecular diffusion. The Batchelor scale [7] is the minimum lengthscale at which scalar fluctuations can persist (Fig. 1), given by the length scale at which fluid deformation balances by molecular diffusion as
$s_{B}=\sqrt{\kappa/\gamma},$ (1)
with $\kappa$ the molecular diffusivity and $\gamma$ the mean stretching rate of the chaotic flow. The value of $\gamma$ is dictated by the accumulation of fluid deformations produced by microscale velocity gradients, scaling as $U/s_{V}$ where $U$ is a characteristic velocity scale. When the characteristic persistence time $T$ of velocity gradients is small compared to the characteristic advection time on the velocity scale $s_{V}/U$, the mean stretching rate is expected [1] to scale as
$\gamma\sim T\,(U/s_{v})^{2}.$ (2)
Flow persistence can be quantified by a dimensionless persistence parameter, or Kubo number [8, 9],
$Ku=\frac{UT}{s_{V}}.$ (3)
For $Ku\ll 1$, the flow has weak persistence while for $Ku\gg 1$, the flow has long time correlations. Eqs.(1) and (2) show that scalar fluctuations are maintained below $s_{v}$ whenever ${\kappa}<U^{2}T$, that is, for weakly diffusing or strongly agitated solutes. In unbounded domains, chaotic flows favour the dispersive spreading of solutes leading to the formation of scalar gradients at increasingly large scales
$s_{d}\sim\sqrt{Dt}$ (4)
where $D$ is the dispersion coefficient and $t$ is the time (Fig. 1). Owing to the Taylor-Green-Kubo relation between velocity fluctuations and dispersion, $D$ generally scales as
$D\sim U^{2}T.$ (5)
Since $s_{d}$ grows in time and $s_{v}$ and $s_{B}$ are constant, we have the general lengthscale separation
$s_{d}\gg s_{v}>s_{B}$ (6)
at least asymptotically in time. Practically, resolving scales smaller than $s_{d}$ or $s_{v}$ is often impossible numerically at small $\kappa$, due to the large number of grid cells necessary to resolve $s_{B}$. Thus, one aims at finding macroscale equations that correctly capture small-scale fluctuations into an effective model. In the majority of chaotic flows, such effective description remains largely unknown.
At the microscale, the lamellar theory [10, 11] accurately predicts the evolution of elementary solute filaments, which elongate and dilute by the coupled action of flow stretching and molecular diffusion. In this one-dimensional representation, the filaments are considered to be independent of each other. Thus, the theory does not allow prediction of mixing statistics in the late time when filaments coalesce and where the concentration field is not Gaussian. Aggregation theories, whether considered above[12] or below [13] the velocity length scale $s_{v}$, have attempted to describe the late-time evolution of the scalar statistics through the evolution of the number of filament coalescence. Well constrained in confined systems [14], aggregation theories are difficult to apply in dispersing interfaces because the density of lamella aggregation becomes a function of both space and time  [15]. For chaotic flows in particular, the exponential increase of material lengths in slowly increasing dispersive volumes (Eq.(4)) questions the applicability of a uniform aggregation process [16], since it would lead to the removal of all scalar fluctuations inside the dispersion zone at a finite time. In contrast, we expect chaotic agitation to maintain fluctuations in proportion to the mean concentration gradients (Fig. 1). In particular, numerical simulations Tsang et al. [17], Haynes and Vanneste [18] in periodic domains have shown that the decay rate of scalar variance is controlled by microscale dissipation processes in domains smaller than a few $s_{v}$, while it is controlled by large-scale gradients on larger domains.
Batchelor [7], Kraichnan [1] and later Kalda [19] proposed an alternative description of mixing by smooth chaotic flow, recognising that concentration gradients are transported to smaller scales by random fluid stirring, while molecular diffusion is only effective at small scales to reduce scalar variance. These authors describe this process through an evolution equation for the power density spectrum (PSD) of scalar fluctuations as a function of time and spatial frequency. While relying on a similar description of the fluid deformation process as the lamellar representation, the spectral description is global in that it is not limited to isolated Gaussian lamellae. Thus, it is well adapted to the case of scalar mixing fronts of arbitrary shape, where single lamellar elements are not distinguishable. To the author’s knowledge, a comparison of such a spectral description with direct numerical simulations of scalar mixing in smooth chaotic flows has never been attempted.
While appealing, the spectral description of mixing does not directly inform on the concentration statistics of the mixture, an information which is critical to many transport applications, e.g. for contaminant spreading or reactive transport. Furthermore, there is currently no consensus on the asymptotic shape of the scalar Probability Density Function (pdf) in smooth chaotic flows, with experimental studies pointing to Gamma [20], Exponential [4], Power-law [17, 21] or even Gaussian pdfs [22]. Theoretical results on the advection-diffusion equation [23] constrain the possible shapes of the scalar pdf, but leave open the underlying physical mechanisms. In particular, the relation between the asymptotic scalar pdf moments of a mixing front and the characteristics of the microscale stirring process has not been not established. In this study, we show that scalar fluctuations within a mixing front are controlled both by the stretching properties of the microscale velocity field and by the dispersive spreading of the scalar. Using Kraichnan [1] model, we propose a physical-based closure for microscale fluctuation and validate it against direct numerical simulations of chaotically agitated fronts.
The paper is organised as follows. In section 2, we recall the theoretical foundations allowing to predict the evolution of scalar variance in mixing fronts and the spectrum of microscale scalar fluctuations. Based on these two elements, we propose a theoretical closure for the scalar dissipation. In section 3, we compare these analytical predictions with direct numerical simulations. We consider a smooth and single scale flow field, the random sine flow, as a prototype of chaotic flow for which the persistence, the dispersivity and the stretching rate can be varied. Sine flows have been extensively used to uncover the key feature of chaotic mixing [24, 25]. Their relative simplicity makes them computationally tractable even for large domains and times. We numerically study the evolution of the concentration fluctuation spectrum in the case of the forcing of large-scale scalar gradient. The numerical spectrum shows the characteristic $k^{-1}$ scaling between $s_{d}$ and $s_{B}$, and accurately matches the Kraichnan’s prediction. We show that a macroscale transport equation coupled with the physically-based closure for chaotic fluctuations accurately capture the evolution of mixing fronts. We finally conclude on the potential of our results to model other transport processes.
II Theory
In this section, we recall earlier results on the behaviour of concentration fluctuations in chaotic flows, notably the evolution of scalar variance and the spectrum of scalar fluctuations. Based on these two elements, we then propose a closure model for microscale fluctuations in the presence of a mean scalar gradient.
II.1 Background
II.1.1 Concentration microscale
The transport of a scalar concentration field $c(\bm{x},t)$ under the action of an incompressible velocity field $\bm{u}(\bm{x},t)$ is governed by the advection diffusion equation
$\partial_{t}c+\bm{u}\cdot\nabla c=\kappa\nabla^{2}c.$ (7)
Taking a Reynolds decomposition of the concentration field between ensemble mean and fluctuations reads
$c(\bm{x},t)$ $=$ $\bar{c}(\bm{x},t)+c^{\prime}(\bm{x},t),$ (8)
$\bm{u}(\bm{x},t)$ $=$ $\bar{\bm{u}}(\bm{x},t)+\bm{u}^{\prime}(\bm{x},t),$ (9)
where $\bar{\bullet}$ is the ensemble mean of the microscale chaotic agitation process, of length-scale $\leq s_{v}$. $\bar{\bullet}$ can also be interpreted as the result of a low-pass filtering with characteristic scale $s_{v}$.
Due to this filtering, we may assume a separation of length and times scales between the mean field $\bar{c}(\mathbf{x})$ and the fluctuation $c^{\prime}(\mathbf{x},t)$, such that the fluctuation satisfies the inhomogeneous ADE
$\partial_{t}c^{\prime}+\bm{u}\cdot\nabla c^{\prime}-\kappa\nabla^{2}c^{\prime}=f(\mathbf{x}),\quad f(\mathbf{x})\equiv-\bm{u}\cdot\nabla\bar{c}-\kappa\nabla^{2}\bar{c},$ (10)
where the source $f$ has zero mean. For temporally-periodic flows with flow period $T$, the fluctuation $c^{\prime}$ in a zero mean field ($\bar{c}=f=0$) is comprised of a finite set of $N$ strange eigemmodes [26, 21, 27], which are Floquet modes of the AD operator $\mathcal{L}$ of the ADE (7) that manifest as exponentially decaying patterns that control the mixing process:
$c^{\prime}(\mathbf{x},t)=\sum_{k=0}^{K}\alpha_{k}(t)\varphi_{k}(\mathbf{x},t)e^{-\lambda_{k}t},$ (11)
where the (possibly complex) eigenmode $\varphi_{k}(\mathbf{x},t)$ is $T$-periodic in time, and the eigenvalue $\lambda_{k}$ (with $|\lambda_{k}|\leqslant|\lambda_{k+1}|$) controls the persistence of the mode. The coefficient $\alpha_{k}$ is governed by the initial condition $c^{\prime}(\mathbf{x},0)$ as is time-dependent due to the non-self adjoint nature of $\mathcal{L}$, where spectral energy is transferred toward the lowest energy (low $k$) modes with time [28]. With time, only the dominant non-trivial mode $\varphi_{0}$ persists, leading to the observation [21, 26] of convergence of $c^{\prime}$ to a persistent yet decaying spatio-temporal pattern in the case of time-period flows. Note strange eigenmode analogues also exist for aperiodic and open flows [29]. In the presence of a source term such as that given by the mean field in (10), the initial conditions for $c^{\prime}$ decay and the fluctuation converges [30] to
[…]
where $p_{k}(t)$ is also $T$-periodic and typically $|p_{k}|\geq|p_{k+1}|$, hence the low-$k$ modes dominate the fluctuation. Hence, the fluctuation $c^{\prime}$ admits truly persistent spatio-temporal patterns due to continual excitation of the natural mixing modes $\varphi_{k}$ by the mean field $\bar{c}$. Equation (12) shows how the transfer of scalar energy between the mean field and small-scale fluctuations shapes the asymptotic structure of the fluctuation and how the natural mixing modes act to dissipate this scalar energy. Although Eqs. (11)–(12) refer to the concentration field in the real space, we will see below that they have direct analogues in the spectral space. The fluctuations $c^{\prime}$ have zero mean and variance
$\sigma_{c}^{2}=\overline{c^{\prime}c^{\prime}},$ (13)
and the mean concentration obeys the isotropic advection dispersion equation at macroscale [31],
$\frac{\partial\bar{c}}{\partial t}+\bar{\bm{u}}\nabla\bar{c}=\nabla\cdot((\kappa+\bm{D})\nabla\bar{c}),$ (14)
with $\bar{\bm{u}}$ the mean flow velocity over the microscale agitation and $\bm{D}$ is the dispersion coefficient obtained with the assumption of a Fickian macroscale flux expression [32]
$\overline{\bm{u}^{\prime}c^{\prime}}\equiv\bm{D}\nabla\bar{c}.$ (15)
When microscale fluctuations are isotropic, and so may be expressed as $\bm{D}=D\mathbf{I}$ The dispersive limit is reached for any velocity fluctuations that have a finite correlation time and integrable statistics [31, 33]. Multiplying Eq. 7 by $c$ and averaging, one obtains an evolution equation for the variance of concentration fluctuations [33, 32]
$\frac{\partial\sigma_{c}^{2}}{\partial t}+\bar{\bm{u}}\nabla\sigma^{2}_{c}-D\nabla^{2}\sigma^{2}_{c}=-2\kappa\overline{(\nabla c^{\prime})^{2}}+2D(\nabla\bar{c})^{2}.$ (16)
The left hand side is the transport of scalar variance through advective and dispersive fluxes while the right-hand side is the dissipation and production of scalar variance due to molecular diffusion and dispersive motion respectively. As shown by Eq.(16), the production of scalar variance is caused by large-scale dispersive fluxes arising from gradients in the mean concentration $\bar{c}$, which bring unmixed fluid patches in contact with each other. In contrast, the destruction of the scalar variance results from microscale diffusive fluxes. Dimensional arguments imply that dissipation can be expressed as
$\overline{(\nabla c^{\prime})^{2}}\sim\sigma^{2}_{c}/s_{m}^{2},$ (17)
where $s_{m}$ is a concentration microscale [34] characteristic of the microscale mixing process. Providing an expression for $s_{m}$ is the main closure problem. Two limiting scenarios can be distinguished depending on the presence of a large-scale forcing process. The first is the decaying scenario where the scalar homogeneizes according to (7), in the absence of any source or boundary conditions. In that case, there is no production of scalar variance and assuming negligible transport,
$\sigma^{2}_{c}\sim\exp{(-2\kappa\,t/s_{m}^{2})},$ (18)
i.e., an exponential decay of scalar variance.The second case arises in the presence of a mean scalar gradient, for instance in a mixing front which is slowly evolving compared to the microscale agitation process. In this case, the left-hand side of Eq. (16) can be neglected, and there is equilibrium between dissipation and production terms. The scalar variance reaches a steady-state with
$\sigma^{2}_{c}=\frac{D}{\kappa}s_{m}^{2}(\nabla\bar{c})^{2}.$ (19)
Large scale concentration gradients spontaneously develop in unbounded mixing domains, due to dispersive effects (Eq. (42)). Thus, the second scenario is the most general and so is considered herein. Since the minimum scale of fluctuations in chaotic flows is fixed on the Batchelor scale, we may first argue that
$s_{m}\sim s_{B}.$ (20)
Inserting this closure in Eq. (19) leads to $\sigma^{2}_{c}=D(\nabla\bar{c})^{2}/\gamma$, independent of $\kappa$. In contrast, numerical simulations [18] generally show a weak dependence on $\kappa$, suggesting that a closure based on the Batchelor scale only is not satisfactory. In the following, we show instead that the transfer of energy from large to small scales globally controls the amount of scalar variance residing in mixing fronts.
II.1.2 Microscale power spectra
As noted by Batchelor [7], the stretching action of the velocity field transfers scalar fluctuations from larger length-scales toward $s_{B}$. The energy spectrum of concentration fluctuations is $E_{k}(k)=|\tilde{c}(k)|^{2}$, where the tilde symbol stands for the spatial Fourier transform
$\tilde{c}=\int c^{\prime}(\bm{x})e^{i\bm{k}\cdot\bm{x}}\text{d}\bm{x}$ (21)
with $k=|\bm{k}|=2\pi/\lambda$ the wavenumber, inversely proportional to the wavelength $\lambda$. The scalar spectrum was shown to present a distinct shape in the two mixing scenarios considered above. In the case of unforced scalar decay, the spectrum is essentially flat above the dissipation scale [17]. In the case of forced mixing (for example, in the presence of a large-scale concentration gradient), the energy spectrum shows a characteristic $E_{k}\sim k^{-1}$ scaling below the injection scale (for $k>2\pi/s_{v}$), suggesting the transfer of large-scale fluctuations to small scales by the stirring action of the flow [7]. Above the injection scale, the spectrum must be essentially Gaussian, reflecting dispersive processes (Eq. (14)). Intermediate cases with $k$ exponents between 0 and -1 were found in numerical simulations with periodic domains with periodicity comparable with $s_{v}$[17, 35] suggesting the transition between micro- and macro-scale mixing mechanisms. The evolution of the scalar spectrum of $c^{\prime}$ was related to the flow stretching statistics by Kraichnan [1] and independently by Kalda [19] in the case of a $\delta$-time correlated random velocity field. The theory rests on the fact that, in the absence of diffusion, the scalar gradients follow the same evolution as the elongation of material elements $\rho=\ell/\ell_{0}$ (or material surfaces in 3D), with $\ell$ and $\ell_{0}$ the final and initial length (or area in 3D) of a fluid element, respectively, which increases due to fluid deformation. The theory is based on three main elements. The first element is that in a $\delta$-correlated flow, the evolution equation for the pdf of elongation $\Lambda(\ln\rho)$ is
$\frac{\partial\Lambda}{\partial t}+\gamma\frac{\partial\Lambda}{\partial\ln\rho}=\sigma^{2}_{\gamma}/2\frac{\partial^{2}\Lambda}{\partial\ln^{2}\rho}.$ (22)
where $\gamma$ is the mean stretching rate and $\sigma^{2}_{\gamma}$ the variance of the stretching rate [19, 36] defined as
$\gamma$ $=$ $\lim_{t\to\infty}\frac{1}{t}\overline{\ln\rho(t)},$ (23)
$\sigma^{2}_{\gamma}$ $=$ $\lim_{t\to\infty}\frac{1}{t}\overline{(\ln\rho(t)-\gamma)^{2}}.$ (24)
Eq. (22) shows that $\Lambda$ evolves according to a mean drift with the mean stretching rate while broadening due to fluctuation in stretching. Note that in $\delta-$correlated chaotic flows, the stretching variability is related to the dimension of the space $d=2,3,\cdots$ by  [1, 25]
$d=2\gamma/\sigma^{2}_{\gamma}.$ (25)
The second element is that an increase in $\ln\rho$ induces the same increment of wavenumber $\ln k$ in the spectral concentration gradient domain, because fluid stretching is compensated for by transverse compression, which transfers scalar gradients to smaller wavelengths. Thus, in absence of molecular diffusion, the power spectra density of scalar gradients $\Phi(k)\sim E_{k}k$ must follow Eq. (22), with $\ln\rho$ replaced by $\ln k$.
The last element is that molecular diffusion becomes efficient at large $k$ to dissipate energy at a rate $2\kappa k^{2}$, which can be included in the evolution equation as
$\frac{\partial\Phi}{\partial t}+\gamma\frac{\partial\Phi}{\partial\ln k}=\frac{\sigma^{2}_{\gamma}}{2}\frac{\partial^{2}\Phi}{\partial\ln^{2}k}-2\kappa k^{2}\Phi.$ (26)
In essence, such model resembles the lamellar description of mixing [37] with the difference that diffusion now operates on an arbitrary concentration field rather than single Gaussian lamellae. This equation has been studied by Kraichnan [1] and Kalda [19]. In the case of a source of large-scale concentration gradients, the boundary condition is $\Phi(0,t)\equiv\chi_{0}/\gamma$, where $\chi_{0}$ is the rate at which scalar variance is injected into the system from these large-scale gradients and $\gamma$ is the mean stretching rate of an infinitesimal line element [7]. Kalda [19] provides an explicit solution to Eq. (26) in terms of Hankel functions. Kraichnan [1] proposes a simpler equivalent form (5.14 in the original paper)
$E_{k}(k)=\frac{\chi_{0}}{\gamma k}f(2ks_{B}^{\prime})\exp(-2ks_{B}^{\prime}).$ (27)
where the length scale
$s_{B}^{\prime}=\sqrt{\kappa/\sigma^{2}_{\gamma}}.$ (28)
is a modified Batchelor scale, and $f(k)$ is the solution of the Kummer’s equation [1]
$kf^{\prime\prime}-(2k+2\gamma/\sigma^{2}_{\gamma}-1)f^{\prime}+(2\gamma/\sigma^{2}_{\gamma}-1)f=0,\quad f(0)=1.$ (29)
As shown by Kraichnan [1], in two dimensions, $\gamma/\sigma^{2}_{\gamma}=1$, and $f\propto k^{1/2}$ at large $k$.
Equation (27) describes scalar fluctuations in the Batchelor regime, that is, when $\gamma$ is not scale-dependant, and shows that for low wavenumbers the energy decays as $k^{-1}$, with an exponential cutoff at high wavenumbers. In contrast, above the characteristic velocity length scale, the mean stretching rate drops and the flow agitation becomes essentially dispersive (Eq. (16)). In the following, we show that the two descriptions (chaotic and dispersive) can be matched on a particular injection scale $s_{i}$, scale at which the scalar variance is transmitted at higher wavenumbers at a rate $\chi_{0}$ controlled by dispersion.
II.2 Proposed microscale closure
As shown in Eq.(16), scalar fluctuations are produced at large scale by the mean dispersive flux. We define the injection scale $s_{i}$ as the characteristic scale at which scalar fluctuations are produced inside the mixing front. In turn, once produced by dispersive motions, scalar fluctuations are transferred to smaller scales by the cascade process described by Kraichnan [1](Eq. (27)). The amount of scalar variance residing in the system between the injection and the dissipation scale (Eq. (13)) can be obtained by integration, e.g.
$\sigma^{2}_{c}=\frac{1}{\pi}\int_{2\pi/s_{i}}^{\infty}E_{k}(k)\,\text{d}k,$ (30)
where $E_{k}$ is given by Eq. (27). The integral is analytical when $f=1$, and gives
$\sigma^{2}_{c}\approx\frac{\chi_{0}}{\pi\gamma}|Ei(-4\pi s_{B}^{\prime}/s_{i})|,$ (31)
where $Ei$ is the exponential integral. We show in Fig. 6 that Eq. (31) accurately describes numerical simulations, even in the case of two-dimensional flows where $f\neq 1$. This is explained by the primary importance of the exponential cutoff in setting the value of the integral (34), rather than the exact shape at intermediate wavenumbers.
The value of $\chi_{0}$ is controlled by the large-scale dispersion process. At large scale, the variance production rate due to mean scalar gradients (Eq. (16)) is:
$\frac{\text{d}}{\text{d}t}\sigma^{2}_{c}=2D(\nabla\bar{c})^{2}.$ (32)
Assuming that this is entirely transmitted to smaller scales, then
$\chi_{0}\equiv\pi\frac{\text{d}}{\text{d}t}\sigma^{2}_{c}=2\pi D(\nabla\bar{c})^{2}.$ (33)
Combining this argument with Eq. (31) provides a complete closure for microscale scalar variance
$\sigma_{c}^{2}\approx\frac{2D(\nabla\bar{c})^{2}}{\gamma}\left|Ei(-4\pi s_{B}^{\prime}/s_{i})\right|.$ (34)
Eq. (34) indicates that the variance of scalar fluctuations in the presence of a mean scalar gradient weakly depends on the strength of molecular diffusion, since the exponential integral $Ei(x)\sim-\ln x$ when $x<1$.
Alternatively, the closure can be formulated on the concentration microscale (Eq. (19)) is
$s_{m}^{2}\approx 2s_{B}^{2}|Ei(-4\pi s_{B}^{\prime}/s_{i})|.$ (35)
Hence, Eq. (35) appears as a correction factor to the simple anzatz $s_{m}\sim s_{B}$ (Eq. (20)). $s_{m}$ can be larger or lower than $s_{B}$ depending on the ratio of injection versus dissipation scales $s_{i}/s_{B}^{\prime}$. Numerical inspection of Eq. (19) shows that $s_{m}>s_{B}$ when $s_{i}/s_{B}^{\prime}>22.7$, and reversely. Since this ratio can be interpreted as a root-squared Péclet number ($Pe\equiv(s_{B}^{\prime}/s_{i})^{2}$), the critical microscale where $s_{m}\approx s_{B}$ is found around $Pe\approx 515$.
Precising the value of $s_{i}$ is the final ingredient of the microsscale closure problem. We envision $s_{i}$ as the typical scale that divides the dispersive scales above $s_{v}$ from the deformation scales below $s_{v}$. It can be estimated as the scale at which the characteristic time to disperse $(s_{i}^{2}/D)$ is of the order of the characteristic time to deform $(\gamma^{-1})$ :
$s_{i}\sim\sqrt{D/\gamma}\propto s_{V},$ (36)
where we have used Eqs. (2) and (5). Note that $s_{i}$ has the same form as the Batchelor scale (Eq. (1)), but with the diffusion coefficient replaced by the dispersion coefficient. Note that this scale has previously been reported to be important in scalar decay through periodic domains, with mixing rates controlled either by local fluid deformation or by global dispersion, respectively below and above a critical domain size, estimated between $2$ and $3s_{v}$ [17, 38]. Recall finally that the theoretical prediction for the fluctuation variance (or equivalently the fluctuation microscale) relies on the following assumptions: (i) the flow is $\delta$ correlated in time, (ii) $\gamma$ is constant (advection is chaotic and scale independent) and (iii) $s_{B}\ll s_{i}$.
In the following section, we numerically test the accuracy of this prediction in a periodic, two-dimensional, smooth, and chaotic flow, the random sine flow. Transport is solved for a fixed mean gradient in a periodic domain (Sections  III.4 and III.5) and in the case of a dispersing front in an infinite domain (Section III.6).
III Comparison to direct numerical simulations
III.1 Methods
To numerically test Eq.(34), we first consider advection-diffusion of passive scalar in the presence of a fixed large-scale gradient $\nabla\bar{c}\equiv\bm{g}$. As done for Eq.(8), the scalar concentration may be split into mean field and fluctuations
$c(x,t)=\bar{c}+c^{\prime}=\bm{g}\cdot\bm{x}+c^{\prime}(x,t)$ (37)
Inserting (37) into the advection diffusion equation (7) leads to
$\partial_{t}c^{\prime}+\bm{u}\cdot\nabla c^{\prime}=\kappa\nabla^{2}c^{\prime}-\bm{u}\cdot\bm{g}$ (38)
As mentioned above, the presence of a source term in the advection-diffusion equation creates a forcing for scalar fluctuations on the scale of $\bm{u}$. Setting $g=0$ recovers the unforced transport equation (7). Without loss of generality, we focus on a mixing front orientated in the $x$ direction, for which the forced gradient is
$\bm{g}=g\,\bm{e}_{x},$ (39)
thus leading to a source term equal to $-gu(y)\bm{e}_{x}$.
III.2 Velocity field
We consider a generic smooth chaotic velocity field, the two-dimensional Zeldovich sine flow [26], or random wave flow. The sine flow is a continuous transformation operating on a periodic domain $[0,s_{v}]\times[0,s_{v}]$, which is defined as $\bm{u}=(u,v)$, with
$v(x,t)$ $=$ $\left\{\begin{array}[]{ll}U\sin(2\pi x/s_{v}+\phi_{j})&\text{ for}jT<t<(j+1/2)T,\\ 0&\text{ for }(j+1/2)T<t<(j+1)T\\ \end{array}\right.,$
$u(y,t)$ $=$ $\left\{\begin{array}[]{ll}0&\text{ for }jT<t<(j+1/2)T,\\ U\sin(2\pi y/s_{v}+\psi_{j})&\text{ for }(j+1/2)T<t<(j+1)T\\ \end{array}\right.,$
with $T$ the time period of the flow, $j$ the time period number and $U$ a constant velocity magnitude. $\phi_{i},\psi_{i}\in[0,\pi]$ are random phases that change at each time period. The velocity field is divergence free. When $Ku\to 0$ (Eq. (3)), the flow changes rapidly over time, causing minimal fluid deformations, while for $Ku\to\infty$ the persistence of the flow strongly affects the deformation of the fluid elements, which tend to align in the main shear directions. The limit $Ku\to 0$ is the $\delta$-time correlated flow envisioned in Kraichnan [1], where analytical results about exist (Eq. (27)). Without loss of generality, we set $T=1$ and $s_{v}=1$ such that the unique free parameter of the flow is $U=Ku$, which controls both the stretching and dispersion statistics of the flow and its persistence. Particle dispersion in random sine flow is given in the long time limit by [17]
$D\equiv\langle\Delta x^{2}+\Delta y^{2}\rangle/(2T)={U^{2}T}/{16}$ (42)
where $\Delta x$ and $\Delta y$ are the mean squared displacement caused by a single flow period. In the Kraichnan limit, the mean stretching rate is given by [25, 1]
$\gamma={\pi^{2}U^{2}T}/({8\,s_{v}^{2}}),$ (43)
while the variance of stretching rates is equal to the mean, e.g. $\sigma_{\gamma}^{2}=\gamma$. Note that $\gamma\to 0$ as $U\to 0$, e.g. in the $\delta$-correlated flow limit. In the case of persistent flows ($UT/s_{V}\gg 0$), material lines partially align with shear directions and such that stretching is less efficient and $\gamma<{\pi^{2}U^{2}T}/({8s_{v}^{2}})$. In addition, the alignement of all material interfaces on the same directions tends to reduce stretching heterogeneity ($\sigma^{2}_{\lambda}<\lambda$). More details about this persistent regime can be found in Meunier and Villermaux [36]. In most random chaotic flows, the elongation of material lines in sine flows approximately follows a log-normal distribution with parameters $\gamma t$ and $\sigma^{2}_{\gamma}t$ that depend on the amplitude $U$, such that Eq. (22) is valid.
III.3 Numerical scheme and convergence
The advection diffusion equation (10) is solved using spectral decomposition in Fourier space. We have checked that the scalar field remains periodic even for $g>0$, such that Eq. (38) can be solved on the unit square.
Following Meunier and Villermaux [25], unidirectional advection with source is solved exactly with Fourier transform in a single coordinate, while diffusion is solved in the full spectral domain. Denoting $\tilde{c}_{x}$ (respectively $\tilde{c}_{y}$) the Fourier transform of $c$ in direction $x$ (respectively $y$), and $\Delta t\leq T/2$ the numerical time step, the advection part in the first half-time period ($iT<t<(i+1/2)T$),
$\frac{\partial\tilde{c}_{x}}{\partial t}+ik_{x}u(y)\tilde{c}_{x}=\mathcal{F}_{x}(-g\,u(y))=-g\,\delta(k_{x})u(y).$ (44)
This yields
$\tilde{c}_{x}(k_{x},y,t+\Delta t)=\tilde{c}_{x}(k_{x},y,t)\exp(-ik_{x}u(y)\Delta t)-g\delta(k_{x})u(y)\Delta t.$ (45)
Diffusion is solved sequentially in the full frequency domain,
$\tilde{c}(k_{x},k_{y},t+dt)=\tilde{c}_{x}(k_{x},k_{y},t)\exp(-\kappa k^{2}\Delta t).$ (46)
The second half period ($(i+1/2)T<t<(i+1)T$) is solved similarly, this time without source term, with
$\tilde{c}_{y}(x,k_{y},t+dt)=\tilde{c}_{y}(x,k_{y},t)\exp(-ik_{y}v(x)\Delta t)$ (47)
[…]
We run simulations for a time large enough to reach stationarity (in a statistical sense). This is achieved for $t\gamma\gtrapprox 10$ when the scalar variance oscillates around a mean constant value (Figure 2b). Significant fluctuation remains around the mean behaviour, with a given correlation time scale that depends on the inverse of the mean stretching rate $\gamma^{-1}$ (Figure 2a). This can be explained by the time a scalar fluctuation is transported from large to small scales, at the mean rate $\gamma$.
Figure 2: a) Time evolution of the scalar variance b) Time convergence of the mean $<\sigma^{2}_{c}>$ over 100 realizations of the random phases for various $U$ at fixed $s_{B}^{\prime}=5.3\cdot 10^{-3}$.
After this transient, the simulated scalar field reached a statistically steady state, previously reported as strange eigenmode [39] or persistent pattern [21]. In fact, the asymptotic state of the forced problem is not an eigenmode itself but rather is given by a combination of Floquet modes (Eq. (12)) excited by the source term. The spatial structure of this asymptotic mixing state appears different in the presence or absence of large-scale forcing, as shown in Figure 3, with large-scale gradients clearly appearing in the forced case. This suggests a different repartition of the modes energy in freely decaying or forced scalar transport. In turn, the persistence of the flow, characterised by the value of $UT/s_{v}$, does not strongly affect the heterogeneity of the concentration field (Fig. 3), although it clearly impacts the alignment of material lines with respect to the (vertical) velocity field. Statistical averages are then calculated over time periods of $40\gamma$ and presented below.
III.4 Spectrum at equilibrium
Figure 3: Concentration $c^{\prime}$ for various amplitude $U$ for forced mixing $g=1$ (top and scalar decay $g=0$ (bottom). In the scalar decay case, concentrations are rescaled by the standard deviation.
Figure 4: Comparison between power density spectrum obtained via numerical simulation (dots) and Eq. (27) (lines) at fixed Péclet number for $U=0.1,0.2$ and $0.8$ (subplots a.1) and a.2)) and at fixed $U=Ku$ and various Péclet numbers (subplots b.1) and b.2)). The columns show the same spectrum in log-log and log-lin scale. We used $\gamma/\chi_{0}=\pi$ to normalise $E_{k}$. Note that we normalized $k$ by $1/s_{B}^{\prime}$ in the right column to evidence the universal exponential cutoff.
We plot the spectrum of $c^{\prime}$ for various flow persistence and Péclet numbers in Figure 4. The numerical spectrum clearly shows the $k^{-1}$ scaling at large wavenumber (Figs. 4a.1 and b.1) and the exponential cutoff at small wavenumbers wavenumber (Figs. 4a.2 and b.2) . We test the accuracy of the Kraichnan model (Eq. (27)) combined with the proposed injection mechanism (Eq.(33)) in predicting the numerical spectrum. Given the theoretical stretching statistics and dispersivity of the sine flow (Eqs. (43) and (42)), and the proposed relationship (Eq. (33)), the scalar variance injection rate is
${\chi_{0}}=\gamma\,\frac{g^{2}}{\pi}.$ (48)
[…]
In the case of persistent flows, the theory accurately captures the spectrum in the small-wavenumber limit but fails to describe the high-wavenumber tails, which significantly depart from the exponential prediction. Flow persistence ($U>0$) results in a steeper cutoff than predicted theoretically (Fig. 4 a.2). This may be attributed to the alignment of the material lines in the main shear direction, which tends to limit the heterogeneity of the exponential stretching and push towards a Gaussian cutoff of the spectrum [7]. We will see later that this is, however, not critical for the prediction of scalar fluctuations.
III.5 Scalar fluctuations at equilibrium
In Fig. 5, we plot the scalar fluctuation pdf as a function of flow persistence and Péclet number. Due to the large-scale forcing, the pdf of fluctuation is stationary in time. The pdf shows a characteristic Gaussian shape, as previously observed in forced mixing cases [23, 22]. The Gaussian shape is preserved in all flow persistence and Péclet numbers. Note also that the pdf of scalar fluctuation without large-scale forcing ($g=0$) have thicker tails than the Gaussian (Fig. 5a), which resemble a power law. The Gaussianity of scalar fluctuations in mixing fronts is an important feature because it allows one to accurately capture the distribution of fluctuations by the second moments only, that is, the scalar variance.
Figure 5: Pdf of $c^{\prime}$ as a function of $U$ (Pe=$10^{5}$) and Péclet ($U=0.4$). The black dotted line correspond to the case without large scale forcing, with the concentration normalized by $\sigma_{c}$. Pdf are ensemble averaged over 100 realisations.
In Fig. 6, we report the value of the scalar variance in several flow and transport conditions and compare it with the theoretical closure proposed herein (Eq. (6)). A best fit of the closure (38) to the numerical data gives
$s_{i}\approx 2.7s_{v},$ (49)
for all flow persistence (Fig. 6). This value is close to the critical domain size that divides globally controlled from locally controlled variance decay rates in sine flows (between 2 and $3s_{v}$  [17, 38]), confirming the validity of the proposed scaling Eq. (36). Note that the value of $s_{i}$ may differ in other flows than the sine flow.
Even if Eq. (27) relies on the assumption of $\delta$-correlated flows, we found that the prediction for scalar variance is excellent even in the case of correlations. This is due to the importance of the universal range $k^{-1}$ in determining the integral (16) when $s_{B}^{\prime}\ll s_{i}$. The divergence between curves at large $s_{B}^{\prime}$ (small $(s_{i}/s_{B}^{\prime})^{2}$) is explained by the increasing role of the cutoff shape in the integral, which departs from exponential at finite persistence time (Fig. 4).
Figure 6: Dependence of the variance of scalar fluctuations $\sigma^{2}_{c}$ upon modified Batchelor scale $s_{B}^{\prime}$ (Eq. (28)) and flow amplitudes ($U$) in the sine flow (colored lines with markers). The theoretical approximation Eq. (31) is plotted in thick dashed line while the exact numerical integration of Eq. (34) is plotted as a thin continuous line. In the sine flow, $\chi_{0}/\gamma=1/\pi$ (Eq. (48)) The injection scale $s_{i}$ is fixed to $s_{i}=2.7s_{v}$, to match the numerical data.
It is striking that even at large Péclet numbers, microscale fluctuations remain limited compared to macroscale scalar gradients. For example, at $Pe=1000$, scalar fluctuations are expected to be less than 10% of the macroscopic gradient. At $Pe=10^{8}$, the fluctuations hardly reach 60% of the macroscopic gradient.
III.6 Evolution of a mixing front
Now, we test the capacity of the proposed closure Eq. (35) to capture the evolution of a unidirectional mixing front, agitated by a chaotic microscale flow (the sine flow), and that spreads to large scales due to dispersion. Since $\bar{c}$ follows the macrodispersion equation with $\overline{\bm{u}}=0$, its solution for an initial sharp step function is
$\bar{c}=(1+\text{erf}(x/\sqrt{4Dt})/2.$ (50)
The evolution of the local mean concentration is well captured by the dispersive equation (Fig. 7a)
Figure 7: Direct numerical simulation of the evolution the moments of a scalar front mixed by a chaotic velocity field (continuous line). (a) Mean concentration and (b) Variance of the concentration (b). The velocity field is the sine flow with parameters $U=0.8$ and $s_{v}=T=1$. Moments are obtained by solving Eq. (7) with $\kappa=7.7\cdot 10^{-7}$, and ensemble averaging over 20 random phases realizations. The theoretical prediction using the closure Eq. (35) is shown in dashed lines.
Using the closure model (19) based on flow stretching statistics, we can deduce the spatio-temporal evolution of scalar variance. Since
$\frac{\partial\bar{c}}{\partial x}=\frac{1}{\sqrt{4\pi Dt}}\exp(-x^{2}/(4Dt)),$ (51)
we have
$\sigma^{2}_{c}(x,t)=\frac{\exp(-x^{2}/(2Dt))}{2\pi t}\left|Ei(-4\pi s_{B}^{\prime}/s_{i})\right|.$ (52)
Eq. (52) accurately capture the evolution of scalar variance along the front at late time (Fig. 7b). At early times ($t\leq 5$), the prediction underestimates the spatial extent of the variance. This is probably due to the non-negligible transport of scalar variance at early time (the left hand side of Eq. (16)), which redistribute fluctuations.
IV Conclusion
We investigated the statistical behaviour of mixing fronts in smooth chaotic flows in the presence of large-scale scalar gradients.
We use the spectral description of mixing in time-uncorrelated flows of Kraichnan [1] to propose a new closure model relating the stretching statistics (mean and variance of the stretching rates) to the local variability of solute concentrations (the scalar variance), depending on the strength of large-scale scalar gradient and the Péclet number. The validity of the approach relies on three key properties of the flow. The first is the smoothness of the flow field, which ensures that stretching is a continuous, scale-independent process below the velocity-length scale. The second one is the scale separation induced by chaotic advection, which implies that the dissipation and production scales of the scalar variance are well separated, the Batchelor scale being well below the characteristic velocity length scale. The third is the rapid time-decorrelation of flow, which ensures that the evolution of stretching statistics can be envisioned as a Fokker-Planck equation (43). We show that this property is, however, not critical, and that a finite amount of flow persistence does not strongly affect the accuracy of the model. In addition, recent results [40] show that $\rho$ follows an Orstein-Uhlenbeck process for steady and transient flows, reducing to Eq. (22) in the Fickian case.
The variance of scalar fluctuations created by a chaotic flow in a mean scalar gradient is a universal function of a Péclet number (Fig. 6), defined as $Pe=\sigma^{2}_{\gamma}s_{i}^{2}/\kappa$, where $\sigma^{2}_{\gamma}$ is the variance of the stretching rates in the chaotic flow (typically of the same order of magnitude of the Lyapunov exponent $\gamma$), $s_{i}$ is a characteristic scale at which scalar variance is produced (typically of the order of the velocity length scale) and $\kappa$ the molecular diffusivity.
We show that the proposed closure captures direct numerical simulations of advection-diffusion transport in a prototype chaotic flow, the sine flow, and is able to predict the evolution of mean and fluctuation of a solute mixing front. Numerical observation also suggests that local concentration statistics are essentially Gaussian, such that the modelling of the mean and variance of concentrations is sufficient to describe the behaviour of scalar plumes in smooth chaotic flows.
The microscopic closure (35) is expected to apply in any smooth single-scale chaotic flow of low persistence. This is the case for steady laminar flows, for instance in Stokes flow through 3D porous media, where the velocity fluctuations induce chaotic advection in the 2D plane transverse to mean flow direction [41]. In such flows, fluid deformations occur on a time scale that is related to the velocity length scale and magnitude ($T\sim s_{v}/U$), such that Eq. (24) simplifies to $\gamma\sim U/s_{v}$. Thus, flow persistence is of order $Ku\sim 1$. In fact, the microscale $s_{m}$ has already been studied in the context of 3D heterogeneous flows at Darcy scale  Kapoor and Gelhar [42]. However, above pore-scale, advection is in general not chaotic [43], and $s_{m}$ is a non-trivial function of time and space, which strongly dependent on initial conditions. In contrast, below the Darcy scale, advection is chaotic and the microscale is a constant, independent of the solute plume geometry. Thus, the closure (35) is expected to be relevant in the context of up-scaling transport from pore to Darcy scale in porous media.
More work is still required to estimate the value of $s_{i}$ in flows other than the sine flow and to confirm its physical origin and universality (Eq. (36)). Numerical test of the closure in 3D-time dependent chaotic flows is also of direct interest. Because it describes the amount of mixing at microscale, the proposed closure opens a new avenue to model non-conservative transport processes, such as reactive processes. Extension to bimolecular reactive transport system can be obtained by focusing on the concentration statistics of two reactive species, with fluctuation product $\overline{c_{A}^{\prime}c_{B}^{\prime}}$ rather than a single one $\overline{c^{\prime}c^{\prime}}$ as done here [44]. This will be the object of a future study.